OpenStudy (mellamokatie):

Write y = (-3/4)x +3 in standard form using integers. A. 3x + 4y = 12 B. 3x + 4y = -12 C. 3x - 4y = 12 D. -3x + 4y =12

OpenStudy (mellamokatie):

@sweetburger

OpenStudy (comrad):

The answer is a letter of the alphabet!

OpenStudy (comrad):

The answer is a letter of the alphabet that falls between A and D!

OpenStudy (sweetburger):

So standard form is in the form Ax+By=C You can not have fractions. Your first step is to move the x value to the left side. You do this by adding 3/4x to both sides.

OpenStudy (mellamokatie):

Oh okay

OpenStudy (sweetburger):

Now your equation will look like this 3/4x+y = 3 Now because we cant have fractions we must clear the denominator by multiplying everything by 4. \[(4)\times \frac{ 3x }{ 4 } + (4)\times y= 3 \times( 4)\]

OpenStudy (sweetburger):

Now everything is in the proper form and you are left with \[3x + 4y = 12\]

OpenStudy (mellamokatie):

Ohh okay that makes sense so then it A

OpenStudy (sweetburger):

yep :)

OpenStudy (mellamokatie):

Awesome!! :D